# Armijo Line-search Makes (Stochastic) Gradient Descent Go Fast

### v1

@@ C.1 Proofs for the Polyak Step-size
We analyze the convergence of GD with the step-size in Eq. 17 under Assumption 3 with $\omega=0$ i.e. we will assume that $f$ is $L$ uniform smooth and that for all $\theta$, $\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)$. From Propositions 1, 5 and 4, we know that this property is true from binary classification with the logistic loss, as well as for multi-class classification with the cross-entropy loss.
@@
For a fixed $\epsilon>0$, assuming $f(\theta)$ is convex, $L$-uniform smooth and satisfies Assumption 3 with $\omega=0$, GD with the Polyak step-size in Eq. 17 and $c>\frac{1}{2}$, requires
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$, where $u$ is an arbitrary comparator s.t. $f(u)<f(\theta_{T})$.
@@
Following a proof similar to that for Theorem 2, for an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
@@
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-f(u)]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assumption 3)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right)^{2}$ (Using Lemma 1 and defining $C:=\frac{(2c-1)}{c^{2}\,\nu}$)
@@
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\gamma:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\gamma$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\gamma}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\gamma}\right)^{2}$
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$ (Using the definition of $\gamma$)
Using the definition of $C$, we conclude that it is sufficient to set $T$ as:
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
@@
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-f(\beta u^{*})}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Theorem 4. In particular, for a comparator $u=\beta u^{*}$, GD with the Polyak step-size and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{(2c-1)}{c^{2}\,\nu}=\frac{(2c-1)}{c^{2}}$,
@@
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{4}\right]$
@@
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if $\alpha:=\frac{f(\beta u^{*})}{\epsilon}$ and if for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon>0$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-\epsilon}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
@@
The first part of the proof considers a general convex $f$ which is $L$ uniform-smooth and satisfies Assumption 3 with $\omega=0$. Assume $\alpha\in(0,1)$ and consider an arbitrary comparator $u$ s.t. $f(u)\leq f(\theta_{T})$ and $\alpha\epsilon\leq f(u)\leq\epsilon$. We set $\eta_{t}=\frac{f(\theta_{t})-\epsilon}{c\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$.
@@
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assumption 3)
@@
Assuming $T$ is the first iterate s.t. $f(\theta_{T})-f(u)\leq\epsilon$, and therefore $f(\theta_{t})-f(u)\geq\epsilon$ that implies $f(\theta_{t})\geq\epsilon+f(u)\geq(1+\alpha)\epsilon$. Therefore
@@
Since $f$ is $L$-uniform smooth, we know that $f(\theta_{T})-f(u)\leq\frac{L}{2}\left\|\theta_{T}-u\right\|_{2}^{2}$. Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{2\,\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{2\epsilon}{L}}{C}\,\left[1+\frac{1}{\alpha}\right]^{2}$
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, $\alpha=\frac{f(\beta u^{*})}{\epsilon}\leq 1$ and we can use the above result. In particular, we choose $u=\beta u^{*}$ and $\theta_{0}=0$ ensuring that $\left\|\theta_{0}-u\right\|=\beta$. Plugging this value in Eq. 22, we conclude that in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
@@
Appendix D Proofs for Section 5
D.1 Proofs for Section 5.1
@@ Appendix D Proofs for Section 5
Using Assumption 4, we know that,
@@
Using Theorem 1 with $R=\frac{1}{\mu^{2}}$ and $L_{0}=0$ completes the proof. ∎
D.2 Proofs for Section 5.2
@@
From Assumption 4 with $\zeta=2$, we know that $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\mu(\theta)\,[f(\theta)-f^{*}]$ and that $f^{*}=0$. Combining these relations, we have that for all $t\leq\tau$,
@@ Appendix E Proofs for Section 6
For an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$, using the SGD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla{f_{t}}(\theta_{t})$,
@@
$\implies-\mathbb{E}\left[{\eta_{t}}\mathbb{E}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
@@
For normalized data, s.t. $\left\|x_{i}\right\|=1$, the logistic regression loss is convex, is uniform smooth with $L=\lambda_{\max}[X^{T}X]=1$, satisfies Assumption 2 with $L_{1}=1$.
@@
For the $T$ defined in the theorem statement, consider two cases:
Case (I): $\mathbb{E}[f(\theta_{T})]<f(u)\leq\epsilon$. This gives the desired result immediately.
Case (II): $\mathbb{E}[f(\theta_{T})]>f(u)$. In this case, we can use the result in Lemma 3. In particular, for the comparator $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon^{2}}\right)$, SGD with stochastic Armijo line-search with $c$, and $\theta_{0}=0$ ensures that if $T$ is the first iteration s.t. $\mathbb{E}[f(\theta_{T})]-f(u)\leq\epsilon\implies\mathbb{E}[f(\theta_{T})]\leq\epsilon\,(1+\epsilon)$, then, for $C:=\frac{(2c-1)}{c\,\lambda_{1}}$, we conclude that,
$\mathbb{E}[f(\theta_{T})]-f(u)$ $\leq\frac{L}{2}\,\mathbb{E}\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{L}{2}\left[\left\|\theta_{0}-u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,\chi^{2}(u)\,T\right]$
$\implies\mathbb{E}[f(\theta_{T})]$ $\leq f(u)+\frac{L}{2}\left[\left\|u\right\|_{2}^{2}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]+2\eta_{\max}\,f(u)\,T\right]$ (Since $\chi^{2}(u)\leq f(u)$)
Setting $\eta_{\max}=\frac{C}{\epsilon}$ ensures that $\eta_{\max}\geq\frac{C}{f(\theta_{t})}$ for all $t\leq T$. Hence,

### v2

@@ C.1 Proofs for the Polyak Step-size
We analyze the convergence of GD with the step-size in Eq. 22 under Assn. 3 with $\omega=0$ i.e. we will assume that $f$ is $L$ uniform smooth and that for all $\theta$, $\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)$. From Prop. 1, 6 and 5, we know that this property is true from binary classification with the logistic loss, as well as for multi-class classification with the cross-entropy loss.
@@
For an initialization $\theta_{0}$, $\epsilon\in(0,f(\theta_{0}))$ and comparator $u$ s.t. $f(u)\leq\epsilon$, if $f(\theta)$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, GD with the Polyak step-size in Eq. 22 and $c>\frac{1}{2}$, requires
$T$ $\geq\frac{c\,\nu^{2}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
iterations to ensure that $f(\theta_{T})-f(u)\leq\epsilon$.
@@
Following a proof similar to that for Thm. 2, for an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
@@
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-f(u)]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assn. 3)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\right)^{2}$ (Using Lemma 1 and defining $C:=\frac{(2c-1)}{c\,\nu^{2}}$)
@@
Assume $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon$. Hence, $\frac{f(\theta_{T})}{f(u)}\leq\alpha:=1+\frac{\epsilon}{f(u)}$. Hence, for all $t<T$, $f(\theta_{t})-f(u)>\epsilon$ and $\frac{f(\theta_{t})}{f(u)}>\alpha$. Consequently, $\frac{f(\theta_{t})-f(u)}{f(\theta_{t})}\geq 1-\frac{1}{\alpha}$. Combining the above relations,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\alpha}\right)^{2}$
Proceeding in the same manner as the proof of Thm. 2, where $B=\frac{e^{q}-1}{q}$ and $q=L_{1}\,\left\|\theta_{0}-u\right\|$, we obtain that,
$f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\underbrace{\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]}_{:=L}\,\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}$
$\implies f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\left(1-\frac{1}{\alpha}\right)\right]=\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\,\frac{\epsilon}{f(u)+\epsilon}\right]$ (23)
Hence, in order to ensure that $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{c\,\nu^{2}\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{f(u)}{\epsilon}\right]^{2}$
@@
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, a fixed $\epsilon\in(0,f(\theta_{0}))$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-f(\beta u^{*})}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Eq. 23 in Thm. 4 with the comparator $u=\beta u^{*}$ where $f(u)\leq\epsilon$. Hence, GD with the Polyak step-size and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{(2c-1)}{c^{2}\,\nu}=\frac{(2c-1)}{c^{2}}$, $L:=\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]$ where $B=\frac{e^{q}-1}{q}$ and $q=L_{1}\,\left\|\theta_{0}-u\right\|$,
@@
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{4}\right]$ (Since $f(\beta u^{*})\leq\epsilon$ and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$)
@@
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if $\alpha:=\frac{f(\beta u^{*})}{\epsilon}$ and if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, a fixed $\epsilon\in(0,f(\theta_{0}))$,, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-\epsilon}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
@@
The first part of the proof considers a general convex $f$ satisfies Assn. 3 with $\omega=0$. Assume $\alpha\in(0,1)$ and consider a comparator $u$ s.t. $\alpha\epsilon\leq f(u)\leq\epsilon$. We set $\eta_{t}=\frac{f(\theta_{t})-\epsilon}{c\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ and note that for all $t$, $f(\theta_{t})\geq\epsilon$.
@@
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assn. 3 with $\omega=0$)
@@
Assuming $T$ is the first iterate s.t. $f(\theta_{T})-f(u)\leq\epsilon$, and therefore $f(\theta_{t})-f(u)\geq\epsilon$ for all $t<T$. Hence, $f(\theta_{t})\geq f(u)+\epsilon\geq(1+\alpha)\epsilon$. Therefore
@@
Proceeding in the same manner as the proof of Thm. 2, where $B=\frac{e^{q}-1}{q}$ and $q=L_{1}\,\left\|\theta_{0}-u\right\|$, we obtain that,
$f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\underbrace{\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]}_{:=L}\,\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}$
$\implies f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\left(1-\frac{1}{(1+\alpha)}\right)^{2}\right]$ (28)
Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\epsilon}{L}}{C}\,\left[1+\frac{1}{\alpha}\right]^{2}$
@@
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, $\alpha=\frac{f(\beta u^{*})}{\epsilon}\leq 1$ and we can use the above result. In particular, we choose $u=\beta u^{*}$ and $\theta_{0}=0$ ensuring that $\left\|\theta_{0}-u\right\|=\beta$ and $f(u)\leq\epsilon$. Plugging this value in Eq. 29, we conclude that in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
@@
Corollary 8.
Assume $f$ is convex and satisfies Assn. 5, then GD-LS with $\eta_{max}=\infty$ requires
$T=\mathcal{O}\left(\frac{R(L_{c}+L_{g}\sqrt{L_{c}}+L_{g}^{2}\sqrt{L_{c}}+L_{g}L_{c})}{\epsilon}+\,\left(R\,(L_{g}(Lg+\sqrt{L_{c}})^{2}+L_{g}(L_{g}+\sqrt{L_{c}}))\right)\,\frac{f^{*}}{\epsilon}\,\ln\left(\frac{f^{*}}{\epsilon}\right)\right)$
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
Proof.
Since $f$ is convex, we can follow a similar argument as in the proof of Cor. 1, to establish that $\left\|\nabla f(\theta_{t})\right\|_{2}^{2}\geq\frac{[f(\theta_{t})-f^{*}]^{2}}{R}$ where $R=\left\|\theta_{0}-\theta^{*}\right\|_{2}^{2}$. This allows us to apply the result from Thm. 1, to obtain a bound on $T$. Since the bound in Case 2 is bigger than Case 1, we consider the bound in Case 2.
$\implies T$ $=\mathcal{O}\left(\frac{\lambda_{0}\,R}{\epsilon}+\,\left(\frac{R\lambda_{1}}{\epsilon}\right)\,\ln\left(\frac{1}{\epsilon}\right)\right)$
From Thm. 1, we have $\lambda_{0}=\frac{3}{1-c}(L_{0}+L_{1}\omega)$ and $\lambda_{1}=\frac{3}{1-c}(L_{1}(\nu+1))$. By using Prop. 3, we can substitute $L_{0},L_{1},\nu,$ and $\omega$ with their respective expressions in terms of $L_{c}$ and $L_{g}$, yielding
$T=\mathcal{O}\left(\frac{R(L_{c}+L_{g}\sqrt{L_{c}}+L_{g}^{2}\sqrt{L_{c}}+L_{g}L_{c})}{\epsilon}+\,\left(R\,(L_{g}(Lg+\sqrt{L_{c}})^{2}+L_{g}(L_{g}+\sqrt{L_{c}}))\right)\,\frac{f^{*}}{\epsilon}\,\ln\left(\frac{f^{*}}{\epsilon}\right)\right)$
C.2 Helper Lemmas
Lemma 5.
For $\epsilon\in(0,M)$ and a comparator $u$ s.t. $f(u)\leq\epsilon$, if $f$ satisfies Assn. 1, 2 and 3 with $L_{0}=0$ and $\omega=0$, then, for all $\theta$ s.t. $\left\|\theta-u\right\|\leq\frac{q}{L_{1}}$,
$f(\theta)-f(u)$ $\leq\frac{\epsilon}{2}+\left[\nu^{2}\,M+B\,L_{1}\,M\right]\frac{\left\|\theta-u\right\|_{2}^{2}}{2}\,,$
where $B:=\frac{e^{q}-1}{q}$.
Furthermore, if $f$ is also $L$ uniform smooth, then, for all $\theta$,
$f(\theta)-f(u)$ $\leq\frac{\epsilon}{2}+\left[\nu^{2}\,M+L\right]\frac{\left\|\theta-u\right\|_{2}^{2}}{2}\,,$
Proof.
Since $f$ satisfies Assn. 2 with $L_{0}=0$, using Eq. 1, we have that for all $\theta$ s.t. $\left\|\theta-u\right\|\leq\frac{q}{L_{1}}$ and $B:=\frac{e^{q}-1}{q}$,
$f(\theta)-f(u)$ $\leq\langle\nabla f(u),\theta-u\rangle+\frac{B\,L_{1}\,f(u)}{2}\left\|\theta-u\right\|_{2}^{2}$
$\leq\frac{\left\|\nabla f(u)\right\|_{2}^{2}}{2\,M\,\nu^{2}}+\frac{\nu^{2}\,M\,\left\|\theta-u\right\|_{2}^{2}}{2}+\frac{B\,L_{1}\,f(u)}{2}\left\|\theta-u\right\|_{2}^{2}$ (Using Young’s inequality)
$\leq\frac{[f(u)]^{2}}{2\,M}+\left[\nu^{2}\,M+B\,L_{1}\,f(u)\right]\frac{\left\|\theta-u\right\|_{2}^{2}}{2}$ (Since $f$ satisfies Assn. 3 with $\omega=0$)
$\leq\frac{\epsilon^{2}}{2\,M}+\left[\nu^{2}\,M+B\,L_{1}\,\epsilon\right]\frac{\left\|\theta-u\right\|_{2}^{2}}{2}$ (Since $f(u)\leq\epsilon$)
$\implies f(\theta)-f(u)$ $\leq\frac{\epsilon}{2}+\left[\nu^{2}\,M+B\,L_{1}\,M\right]\frac{\left\|\theta-u\right\|_{2}^{2}}{2}$ (Since $\frac{\epsilon^{2}}{2\,M}\leq\frac{\epsilon}{2}$ for $\epsilon\leq M$)
If $f$ is $L$-uniform smooth, we can use the standard descent lemma to get, that for all $\theta$,
$f(\theta)-f(u)$ $\leq\langle\nabla f(u),\theta-u\rangle+\frac{L}{2}\left\|\theta-u\right\|_{2}^{2}$
Using this inequality and following the same proof gives the result. ∎
Appendix D Proofs for Sec. 5
D.1 Proofs for Sec. 5.1
@@ C.2 Helper Lemmas
Using Assn. 4, we know that,
@@
Using Thm. 1 with $R=\frac{1}{\mu^{2}}$ and $L_{0}=0$ completes the proof. ∎
See 4
Proof.
From Prop. 4, we know that the MAB problem satisfies both Assn. 2 and Assn. 3 with the parameters $L_{0}=0$, $L_{1}=72$, $\nu=24$, and $\omega=0$. It also satisfies Assn. 4 with $\zeta=1$. Since GD-LS guarantees monotonic decrease of the objective,  Mei et al. (2020, Lemma 5 and Proposition 2), implies that for the uniform initialization where $\pi_{0}(a)=\frac{1}{K}$ for all $a$, $\pi_{\theta_{t}}(a^{*})\geq\frac{1}{K},\,\,\forall\,\,t\geq 0$. Therefore, we have $\mu=\min_{t}\pi_{\theta_{t}}(a^{*})=\frac{1}{K}$. With these parameters, we can instantiate the result of Cor. 3, and get:
[…]
$\implies T$ $=O\left(K^{2}\ln\left(\frac{1}{\epsilon}\right)\right)$ (since $f^{*}=0$ and $f(\theta_{0})-f^{*}\leq 1$)
Corollary 9.
For the tabular MDP problem defined in Prop. 7, GD-LS, with $c=\frac{1}{2}$, and $\eta_{max}=\infty$ requires
$T\geq\frac{2^{7}3}{\mu}\left(\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{3}\,S^{\frac{3}{2}}\right)\ln\left(\frac{1}{(1-\gamma)\epsilon}\right)$
iterations to guarantee $f(\theta_{T})=V^{\pi^{*}}(\rho)-V^{\pi_{\theta_{T}}}(\rho)\leq\epsilon$, where $\mu=\inf_{t\in[T]}\mu(\theta_{t})$ and $\mu(\theta)=\frac{\min_{s}\pi_{\theta}(a^{*}(s)|s)}{\sqrt{S}\,\min_{s\in{\mathcal{S}}}\rho(s)}$, $a^{*}(s)$ is the action corresponding to the optimal policy $\pi^{*}$ in state $s$, $\gamma$ is the discount factor, $\rho$ is the initial state distribution, and $S$ is the size of state space.
Proof.
As proved in Prop. 7, the function $f(\theta)$ satisfies Assn. 2 with $L_{0}=0$ and $L_{1}=8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{2}\,S$,  Assn. 3 with $\nu=8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\sqrt{S}$ and $\omega=0$ and  Assn. 4 with . Since GD-LS guarantees monotonic decrease in the objective,  Mei et al. (2020, Lemma 9), implies that $\mu>0$. With these parameters and noting that $f^{*}=0$, we can instantiate the result of Cor. 3 and get:
$T\geq\max\left\{1,\frac{48}{\mu}\left(\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{2}\,S\left(8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\sqrt{S}+1\right)\right)\right\}\ln\left(\frac{f(\theta_{0})}{\epsilon}\right)$ (since $\lambda_{1}=3\frac{L_{1}(\nu+1)}{1-c}$ and $c=\frac{1}{2}$)
$\implies T\geq\frac{48}{\mu}\left(\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{2}\,S\left(8\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]\,\sqrt{S}+1\right)\right)\ln\left(\frac{1}{(1-\gamma)\epsilon}\right)$ (since $V^{\pi}(\rho)\leq\frac{1}{1-\gamma}$ for all policies $\pi$)
$\implies T\geq\frac{2^{7}3}{\mu}\left(\left[3+\frac{4\cdot\left(\min_{s}\frac{1}{\rho(s)}-(1-\gamma)\right)}{1-\gamma}\right]^{3}\,S^{\frac{3}{2}}\right)\ln\left(\frac{1}{(1-\gamma)\epsilon}\right)$
D.2 Proofs for Sec. 5.2
@@
From Assn. 4 with $\zeta=2$, we know that $\left\|\nabla f(\theta)\right\|_{2}^{2}\geq\mu(\theta)\,[f(\theta)-f^{*}]$ and that $f^{*}=0$. Combining these relations, we have that for all $t\leq\tau$,
@@
Using the SGD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla{f_{t}}(\theta_{t})$, for the comparator $u$ s.t. $f(u)\leq f(\theta_{t})$ for all $t\in[T]$,
@@
$\implies-\mathbb{E}\left[{\eta_{t}}\left[{f_{t}}(\theta_{t})-{f_{t}}^{*}\right]\right]$ $\leq-\min\left\{\eta_{\max},\frac{C^{\prime}}{f(\theta_{t})}\right\}\,\mathbb{E}[f(\theta_{t})-f(u)]$
@@
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$, satisfies Assn. 2 with $L_{0}=0$, $L_{1}=8$ and Assn. 3 with $\nu=8$, $\omega=0$.
@@
For the $\epsilon$ and $T$ defined in the theorem statement, define the event $G=\{\inf_{t\in[T]}[f(\theta_{t})-f(u)]\leq\epsilon\}$. Using the law of total expectation,
$\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]\right]$ $=\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G\right]\,\Pr[G]+\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]\,\Pr[G^{c}]$ (37)
$\leq\epsilon+\mathbb{E}\left[\inf_{t\in[T]}[f(\theta_{t})-f(u)]|G^{c}\right]$ (By definition of $G$)